# A Comprehensive Survey on Synthetic Infrared Image synthesis

### v1

@@ 4.12 Multispectral Pedestrian Detection Dataset [20]
| 1 | OSU Thermal Pedestrian Database | 284 | LWIR | 360x240 | 8 | Raytheon 300D Thermal Sensor |
| 2 | OSU Color-Thermal Database | 17089 | RGB and LWIR | 320x240 | 24bit RGB \& 8bit thermal | Sony TRV87 and Raytheon PalmIR 250D thermal sensor |
| 3 | Terravi Databases | 18 Video Sequences | LWIR | 320x240 | 8 | L-3 Thermal-Eye 2000AS sensor |
| 4 | CSIR-CSIO | 18 Video Sequences | LWIR | 640x480 | ||
| 5 | ICRA ASL-TIR | 4381 | LWIR | 324x256 | 8 | FLIR Tau 320 thermal infrared camera |
| 6 | VAIS | 2865 | RGB and LWIR | 5056x5056 and 1024x68 | ISVI IC-C25 and Sofradir-EC Atom 1024 | |
| 7 | BU-TIV | 65590 | LWIR | upto 1024x | 16 | FLIR SC8000 |
| 8 | CVC-14 | 8385 | RGB and LWIR | 476x640 | ||
| 9 | CVC-09 | 11071 | LWIR | 640x480 | ||
| 10 | FLIR-ADAS | 26000 | RGB and LWIR | 640x512 | 14bit Tiff, 8bit JPEG , 8bit RGB | Teledyne FLIR Blackfly S BFS-U3-51S5C (IMX250) camera and Teledyne FLIR Tau 2 thermal sensor |
| 11 | LLVIP | 30976 | RGB and LWIR | 1920x1080 and 1280x720 | 8 and 16 | HIKVISION DS-2TD8166BJZFY-75H2F/V2 |
| 12 | TNO | 16 Video sequences | RGB, NIR and LWIR | 256x256 | 8 | TRICLOBS (TRI-band Color Low-light OBServation) all-day all-weather surveillance system |
| 13 | LTIR | 11260 | LWIR | 1920x480 | 8 and 16 | FLIRA35, FLIR Tau320, FLIRA655SC |
| 14 | KAIST | 95000 | RGB and LWIR | 320x256 | 8 | FLIR A35 and PointGrey Flea3 |
| 15 | ATR | 600000+ | RGB and MWIR | 640x480 | 16 | Illunis Camera with Nikon Zoom Lens and Night Conqueror MWIR imager |

### v2

@@ 4.12 Multispectral Pedestrian Detection Dataset [20]
1 | OSU Thermal Pedestrian Database | 284 | LWIR | 360x240 | 8 | Raytheon 300D Thermal Sensor |
2 | OSU Color-Thermal Database | 17089 | RGB and LWIR | 320x240 | 24bit RGB \& 8bit thermal | Sony TRV87 and Raytheon PalmIR 250D thermal sensor |
3 | Terravi Databases | 18 Video Sequences | LWIR | 320x240 | 8 | L-3 Thermal-Eye 2000AS sensor |
4 | CSIR-CSIO | 18 Video Sequences | LWIR | 640x480 | ||
5 | ICRA ASL-TIR | 4381 | LWIR | 324x256 | 8 | FLIR Tau 320 thermal infrared camera |
6 | VAIS | 2865 | RGB and LWIR | 5056x5056 and 1024x68 | ISVI IC-C25 and Sofradir-EC Atom 1024 | |
7 | BU-TIV | 65590 | LWIR | upto 1024x | 16 | FLIR SC8000 |
8 | CVC-14 | 8385 | RGB and LWIR | 476x640 | ||
9 | CVC-09 | 11071 | LWIR | 640x480 | ||
10 | FLIR-ADAS | 26000 | RGB and LWIR | 640x512 | 14bit Tiff, 8bit JPEG , 8bit RGB | Teledyne FLIR Blackfly S BFS-U3-51S5C (IMX250) camera and Teledyne FLIR Tau 2 thermal sensor |
11 | LLVIP | 30976 | RGB and LWIR | 1920x1080 and 1280x720 | 8 and 16 | HIKVISION DS-2TD8166BJZFY-75H2F/V2 |
12 | TNO | 16 Video sequences | RGB, NIR and LWIR | 256x256 | 8 | TRICLOBS (TRI-band Color Low-light OBServation) all-day all-weather surveillance system |
13 | LTIR | 11260 | LWIR | 1920x480 | 8 and 16 | FLIRA35, FLIR Tau320, FLIRA655SC |
14 | KAIST | 95000 | RGB and LWIR | 320x256 | 8 | FLIR A35 and PointGrey Flea3 |
15 | ATR | 600000+ | RGB and MWIR | 640x480 | 16 | Illunis Camera with Nikon Zoom Lens and Night Conqueror MWIR imager |